OpenStudy (anonymous):

Choose the correct simplification of (x7)2.

OpenStudy (campbell_st):

if its \[(x^7)^2 = x^{7 \times 2} = x^{14}\]

OpenStudy (anonymous):

okay thats what i thought thanks :D

OpenStudy (anonymous):

then if i have a (7x^3y^3)^2 i would multiply 7 by 2 and then add the others? or multiply them also?

OpenStudy (campbell_st):

this question has several parts as the power of 2 acts on each part \[(7x^3y^3)^2 = 7^2x^{3 \times2}y^{3\times2} = 49x^6y^6\]

OpenStudy (anonymous):

oh okay